OpenStudy (anonymous):

Use the substitution method to solve the system of equations. Enter your answer as an ordered pair 14x - 2y = 78 2x - 2y = 6

OpenStudy (anonymous):

2x-2y=6 divide this with 2 x-y=3 x=3+y..(1) sub to 14x - 2y = 78 14(3+y)-2y=78 42+14y-2y=78 12y=36 y=3 sub to x=3+y..(1) x=3+3=6 (3,6)